Can you please help me with these questions? I am stuck!
1. Simplify completely. 5x - 10/x^2 divided by 5 =
-2,
x - 2/x,
-1, or
x^2 - 2x/x^2
?
2. Simplify the following expression. Please show all work.
x^2-25/x divided by x^2 + 6x + 5/x^2
3. Simplify completely.
4a - 8/6-12a =
-2/3,
2(1-2)/3(1-2a),
2a-2/3-6a,
or 2a-2/3-3a?
4. Simplify completely.
3x + 9/3x =
9,
4,
x+3/x,
or x+6?
5. Simplify completely. x^2 - 25/x+1 multiplied by x^2 - 1/x -5 = (x-5)(x-1),(x+5),(x+1),(x+5)(x-1), or (x-5)(x+1)
6. Choose all the polynomials that are classified as difference of squares.
p^2+4,
9n^2-1,
4a-b^2,
12a^2 - 10a,
m^4-n^4
7. Add and simplify.
x/3+5x/3 =
2x,
5x^2/9,
x,
or 5x^2/3
8. Add or subtract and simplify completely.
A. 16+x/x^3 + 7-4x/x^3
and B. 5-t/t^2 + 1 - t - 3/ t^2+1
9. Simplify completely.
2x+3/x-2 - x+5/x-2 =
x+8/ x- 2,
3x-2/x-2,
1,
x-2/x
10. Solve for x.
Solve x^2/x+12 = 2/1...
-8,3
24, -1
-12,0 or
6,-4
11. Subtract and simplify
x^2/x+3 - 9/x+3 =
x^2-9/x+3,
x - 3,
x^2-9/x^2+9,
or x^2-9/2x+6
12. Subtract and simplify
2y+7/y-5 - y+12/y-5 =
y^2-84/y-5,
1,
y+19/y-5,
or zero

13.solve for x.
x+2/3 = 6/ x -1
-5, 4
16, 19
-2,1
or -4,5
Thank you!!

Answer by edjones(8007)

11)
x^2/(x+3) - 9/(x+3) Use parentheses so tutors will answer your questions.
=(x^2-9)/(x+3)
=(x+3)(x-3)/(x+3)
=x-3
.
Ed